# Tracial states on groupoid $C^{*}$-algebras and essential freeness

### v2

@@ Abstract.
As applications, we obtain the following: 1) a group action is essentially free if every tracial state on the reduced crossed product is canonical and every isotropy group is amenable; 2) if the groupoid $\mathcal{G}$ is second countable, amenable and essentially free then every (not necessarily faithful) tracial state on the reduced groupoid $C^{*}$-algebra is quasidiagonal.
@@
In recent years, there is increasing interest in the classification of crossed products of $C^{*}$-algebras $C_{0}(X)\rtimes_{\nu}\Gamma$ arising from discrete amenable group actions on locally compact Hausdorff spaces $\Gamma\curvearrowright X$ (see, e.g., [2, 10, 12, 15, 16, 17, 19, 22, 24, 25, 28, 29]). One of the key ingredients in those proofs is that every tracial state $\tau$ on $C_{0}(X)\rtimes_{\nu}\Gamma$ is canonical in the sense that $\tau=\tau|_{C_{0}(X)}\circ E$, where $E:C_{0}(X)\rtimes_{\nu}\Gamma\to C_{0}(X)$ is the canonical conditional expectation. Actually, it was shown in [14, Theorem 2.7] that every tracial state on the maximal crossed product $C(X)\rtimes\Gamma$ of an action on a compact Hausdorff space $X$ is canonical if and only if the action is essentially free with respect to all invariant probability Radon measures on $X$. While every tracial state on the reduced crossed product $C(X)\rtimes_{r}\Gamma$ is canonical if and only if the action of the amenable radical $R_{a}(\Gamma)$ of $\Gamma$ (i.e., the largest amenable normal subgroup in $\Gamma$) on $X$ is essentially free with respect to all invariant probability Radon measures on $X$ (see [33, Corollary 1.12]).
On the other hand, X. Li was able to show that all classifiable $C^{*}$-algebras necessarily arise from twisted étale groupoids (see [20]). Hence, it is natural to consider canonical tracial states on general groupoid $C^{*}$-algebras $C_{\nu}^{*}(\mathcal{G})$ of locally compact Hausdorff étale groupoids $\mathcal{G}$. Similarly, a tracial state $\tau$ on $C_{\nu}^{*}(\mathcal{G})$ is canonical if $\tau=\tau|_{C_{0}(\mathcal{G}^{(0)})}\circ E$, where $E:C_{\nu}^{*}(\mathcal{G})\to C_{0}(\mathcal{G}^{(0)})$ is the canonical conditional expectation. If we let $\mu$ be the uniquely associated invariant probability Radon measure on $\mathcal{G}^{(0)}$ to $\tau|_{C_{0}(\mathcal{G}^{(0)})}$, then it follows from [26, Corollary 1.2] and [27, Corollary 2.4] that if $\mathcal{G}$ is second countable, then a tracial state $\tau$ on the maximal groupoid $C^{*}$-algebra $C^{*}(\mathcal{G})$ is canonical if and only if $\mathcal{G}$ is essentially free with respect to the associated measure $\mu$ (see Definition 3.3).
@@
where $f_{i}\in C_{0}(X)$ and $\gamma_{i}\in\Gamma$ for $i=1,\cdots,n$, and $\mathrm{Fix}(\gamma_{i})=\{x\in X:\gamma_{i}x=x\}$ is the set of fixed points of $\gamma_{i}$. The key observation is that $\tau^{\mathrm{Fix}}_{\mu}$ defined on $C_{c}(\mathcal{G})$ is canonical if and only if $\mathcal{G}$ is essentially free with respect to $\mu$ (see Lemma 3.14). As a consequence, it leads to the following main result of this paper, which answers Question 1.1:
@@
(Corollary 5.5) Let $\Gamma$ be a discrete group acting on a locally compact Hausdorff space $X$ equipped with an invariant probability Radon measure $\mu$. We consider the following conditions:
@@
Then (1) normal-$\Rightarrow$ (2) [^1] Note that the implication “(2) normal-$\Rightarrow$ (1)” in Corollary B does not hold in general (see Remark 5.6). and (2) + (3) normal-$\Rightarrow$ (1).
If additionally $\Gamma$ is countable and $X$ is second countable, then (1) normal-$\Leftrightarrow$ (2) + (3).
We end this paper by the following result about quasidiagonal traces, which play a crucial role in the classification of simple nuclear $C^{*}$-algebras (see [35]):
@@
(Corollary 5.15222We also refer the reader to Theorem 5.14 for a more general result.) Let $\mathcal{G}$ be a locally compact, Hausdorff, second countable, amenable and étale groupoid, which is also essentially free. Then every (not necessarily faithful) tracial state on $C_{r}^{*}(\mathcal{G})$ is quasidiagonal.
@@ 2.1. Basic notions for groupoids
and an inverse map on $\mathcal{G}$ given by $\gamma\mapsto\gamma^{-1}$. These operations satisfy a couple of axioms, including the associativity law and the fact that elements in $\mathcal{G}^{(0)}$ act as units. For $x\in\mathcal{G}^{(0)}$, we denote $\mathcal{G}^{x}:=\mathrm{r}^{-1}(x)$ and $\mathcal{G}_{x}:=\mathrm{s}^{-1}(x)$. Moreover, $\mathcal{G}^{x}_{x}=\mathcal{G}^{x}\cap\mathcal{G}_{x}$ is called the isotropy group at $x\in\mathcal{G}^{(0)}$. A subset $Y\subseteq\mathcal{G}^{(0)}$ is called invariant if $\mathrm{r}^{-1}(Y)=\mathrm{s}^{-1}(Y)$, and we denote $\mathcal{G}_{Y}:=\mathrm{s}^{-1}(Y)$. For $A,B\subseteq\mathcal{G}$, we denote
@@
A locally compact Hausdorff groupoid is a groupoid $\mathcal{G}$ endowed with a locally compact and Hausdorff topology for which composition, inversion, source and range maps are continuous with respect to the induced topologies. A subset $A\subseteq\mathcal{G}$ is called a bisection if the restrictions of $\mathrm{s},\mathrm{r}$ to $A$ are homeomorphisms onto their respective images.
We say that a locally compact Hausdorff groupoid $\mathcal{G}$ is étale if the range (and hence the source) map is a local homeomorphism, i.e., for any $\gamma\in\mathcal{G}$ there exists an open neighbourhood $U$ of $\gamma$ such that $\mathrm{r}(U)$ is open and $\mathrm{r}|_{U}$ is a homeomorphism. In this case, the fibers $\mathcal{G}_{x}$ and $\mathcal{G}^{x}$ with the induced topologies are discrete for each $x\in\mathcal{G}^{(0)}$, and $\mathcal{G}^{(0)}$ is clopen in $\mathcal{G}$. Moreover, it follows from definitions that all open bisections form a basis for the topology of $\mathcal{G}$. As a direct consequence, any function $f\in C_{c}(\mathcal{G})$ can be written as a linear combination of continuous functions whose supports are contained in pre-compact open bisections.
We record the following known result (see, e.g., [31, Lemma 8.4.11]). For convenience of the reader, we provide here a short proof.
@@
Given two open subsets $U$ and $V$ of $\mathcal{G}$, we need to show that $UV$ is also open. Without loss of generality, we can assume that $U$ and $V$ are open bisections with $\mathrm{s}(U)=\mathrm{r}(V)$. Fix $\alpha\in U$ and $\beta\in V$ with $\mathrm{s}(\alpha)=\mathrm{r}(\beta)$. Since the multiplication map is continuous at $(\alpha^{-1},\alpha\beta)$, there exists an open neighbourhood $U_{1}\subseteq U$ of $\alpha$ and an open bisection $W\subseteq\mathrm{r}^{-1}(\mathrm{r}(U_{1}))$ containing $\alpha\beta$ such that $(U_{1}^{-1}\times W)\cap\mathcal{G}^{(2)}\subseteq V$. As $U_{1}\subseteq U$ is also a bisection, we obtain that $W\subseteq U\cdot V$. ∎
@@ 2.2. Groupoid $C^{*}$-algebras
Let us now recall different constructions of groupoid $C^{*}$-algebras and their basic properties. Given a locally compact Hausdorff and étale groupoid $\mathcal{G}$, the space $C_{c}(\mathcal{G})$ can be turned into a ***-algebra with the following operations: let $f,g\in C_{c}(\mathcal{G})$, we define their convolution and involution
@@
We could also consider the point-wise multiplication $f\cdot g$ for $f,g\in C_{c}(\mathcal{G})$.
@@
and the reduced groupoid $C^{*}$-algebra $C^{*}_{r}(\mathcal{G})$ is defined to be the completion of the ***-algebra $C_{c}(\mathcal{G})$ with respect to the reduced $C^{*}$-norm $\|\cdot\|_{r}$. It is clear that each left regular representation $\lambda_{x}$ can be extended automatically to a ***-homomorphism $\lambda_{x}:C^{*}_{r}(\mathcal{G})\to\mathscr{B}(\ell^{2}(\mathcal{G}_{x}))$.
@@
We will also consider other $C^{*}$-norms between the reduced and the maximal ones. More precisely, we say that a $C^{*}$-norm $\|\cdot\|_{\nu}$ dominates the reduced $C^{*}$-norm if $\|f\|_{\nu}\geq\|f\|_{r}$ for any $f\in C_{c}(\mathcal{G})$. It is worth noticing that $\|\cdot\|_{\nu}\leq\|\cdot\|_{\max}$ always holds. We denote the $C^{*}$-completion of $C_{c}(\mathcal{G})$ with respect to $\|\cdot\|_{\nu}$ by $C^{*}_{\nu}(\mathcal{G})$, called a groupoid $C^{*}$-algebra of $\mathcal{G}$. Similarly, we have a surjective $\ast$-homomorphism
@@
We remark that there is an inclusion map $\iota_{0}:C_{c}(\mathcal{G}^{(0)})\to C_{c}(\mathcal{G})$ given by extending functions by zero on $\mathcal{G}\setminus\mathcal{G}^{(0)}$, and it was recorded in [3, Section 2.2] that $\iota_{0}$ can be extended to an isometric $\ast$-homomorphism $\iota:C_{0}(\mathcal{G}^{(0)})\hookrightarrow C^{*}_{r}(\mathcal{G})$, where the norm on $C_{0}(\mathcal{G}^{(0)})$ is given by the supremum norm. The same fact holds for any $C^{*}$-norm $\|\cdot\|_{\nu}$ dominating the reduced $C^{*}$-norm. Hence, we will in the sequel regard $C_{0}(\mathcal{G}^{(0)})$ as a $C^{*}$-subalgebra in $C^{*}_{\nu}(\mathcal{G})$ without further explanation.
From [30, Proposition II.4.2] (see also [4, Section 2.2]) we have that any element in $C^{*}_{r}(\mathcal{G})$ can be regarded as a $C_{0}$-function on the groupoid $\mathcal{G}$. Indeed, there exists a linear and contractive map $j:C^{*}_{r}(\mathcal{G})\to C_{0}(\mathcal{G})$ given by
@@
for $a\in C^{*}_{r}(\mathcal{G})$ and $\gamma\in\mathcal{G}$. On $C_{c}(\mathcal{G})\cup C_{0}(\mathcal{G}^{(0)})$ the map $j$ is nothing but an identity. The reduced groupoid $C^{*}$-algebra also admits a faithful conditional expectation $E:C^{*}_{r}(\mathcal{G})\to C_{0}(\mathcal{G}^{(0)})$ defined by
@@
We end this subsection with an elementary fact, and leave its proof to the reader.
@@ 2.3. Tracial states
Given a tracial state $\tau$ on the groupoid $C^{*}$-algebra $C^{*}_{\nu}(\mathcal{G})$, then $\tau|_{C_{0}(\mathcal{G}^{(0)})}$ is a state on $C_{0}(\mathcal{G}^{(0)})$, which corresponds to a (positive) probability Radon measure $\mu$ on $\mathcal{G}^{(0)}$ according to the Riesz representation theorem. In other words, we have
$\tau(f)=\int_{\mathcal{G}^{(0)}}f\mathrm{d}\mu\quad\text{for any}\quad f\in C_{0}(\mathcal{G}^{(0)}).$
Such $\mu$ is called the associated measure to $\tau$ and in this case also denoted by $\mu_{\tau}$. It is actually invariant in the following sense.
@@
(2.6) $\alpha_{B}:\mathrm{s}(B)\to\mathrm{r}(B)\quad\text{by}\quad x\mapsto\mathrm{r}((\mathrm{s}|_{B})^{-1}(x))\quad\text{for}\quad x\in\mathrm{s}(B).$
A Borel measure $\mu$ on $\mathcal{G}^{(0)}$ is called invariant (cf. [30, Definition I.3.12]) if for any open bisection $B$ in $\mathcal{G}$, we have $\mu|_{\mathrm{r}(B)}=(\alpha_{B})_{\ast}(\mu|_{\mathrm{s}(B)})$. We also denote
@@
Conversely, let $\mu$ be an invariant probability Radon measure on $\mathcal{G}^{(0)}$. Then it follows from [21, Lemma 4.2] (see also [30, Proposition II.5.4]) that
@@
is a tracial state on $C^{*}_{r}(\mathcal{G})$, called the tracial state associated to $\mu$. Similarly, we can also consider the associated tracial state to $\mu$ on $C^{*}_{\nu}(\mathcal{G})$ (with the same notation)
@@
Let $\mathcal{G}$ be a locally compact Hausdorff and étale groupoid, and $\|\cdot\|_{\nu}$ be a $C^{*}$-norm on $C_{c}(\mathcal{G})$ dominating the reduced $C^{*}$-norm. A tracial state $\tau$ on $C^{*}_{\nu}(\mathcal{G})$ is called canonical if $\tau=\tau_{\mu}$, where $\mu$ is the associated measure to $\tau$ on $\mathcal{G}^{(0)}$.
The following is elementary but useful in the sequel:
@@
A tracial state $\tau$ on $C^{*}_{\nu}(\mathcal{G})$ is canonical if and only if $\tau(f)=0$ for any $f\in C_{c}(\mathcal{G}\setminus\mathcal{G}^{(0)})$.
@@
One implication is clear. Now we assume that $\tau(f)=0$ for any $f\in C_{c}(\mathcal{G}\setminus\mathcal{G}^{(0)})$. As $C_{c}(\mathcal{G})$ is dense in $C^{*}_{\nu}(\mathcal{G})$, it suffices to show that $\tau(f)=\tau_{\mu}(f)$ for any $f\in C_{c}(\mathcal{G})$, where $\mu$ is the associated measure to $\tau$. We note that étaleness and Hausdorffness of $\mathcal{G}$ imply that $\mathcal{G}^{(0)}$ is clopen in $\mathcal{G}$ so that $\tau(f)=\tau_{\mu}(f)$ for $f\in C_{c}(\mathcal{G}^{(0)})$ by the definition of $\mu$. Finally, we conclude the proof by the assumption. ∎
It is interesting to know when every tracial state on $C^{*}_{\nu}(\mathcal{G})$ is canonical. Recall from [21, Lemma 4.3] (see also [30, Proposition II.5.4]) that if the groupoid $\mathcal{G}$ is principal, then every tracial state on $C^{*}_{r}(\mathcal{G})$ is canonical. We would like to weaken the condition of principalness to the following notion of essential freeness.
@@
We start with the following proposition, which is perhaps known to experts at least for the maximal $C^{*}$-norm (see [26, Corollary 1.2]). We provide here a self-contained proof, because we cannot find the explicit statement we need in the literature.
@@
Our proof here is mainly inspired by the one for [18, Proposition 1.1]. By Lemma 3.2, it suffices to show that $\tau(f)=0$ for any $f\in C_{c}(\mathcal{G}\setminus\mathcal{G}^{(0)})$. Using an argument of partitions of unity, we can additionally assume that $\mathrm{supp}f\subseteq B$ for some pre-compact open bisection $B$ such that $\overline{B}\subseteq B_{0}$ for anther pre-compact open bisection $B_{0}\subseteq\mathcal{G}\setminus\mathcal{G}^{(0)}$.
First we suppose that $\mathrm{supp}f\cap\{\gamma\in B:\mathrm{s}(\gamma)=\mathrm{r}(\gamma)\}=\emptyset$. Then for any $\gamma\in\mathrm{supp}f$, there exists an open neighbourhood $W_{\gamma}$ of $\gamma$ such that $\mathrm{s}(W_{\gamma})\cap\mathrm{r}(W_{\gamma})=\emptyset$. Using an argument of partitions of unity, we can further assume that $\mathrm{supp}f\subseteq W$ for some open $W\subseteq B$ such that $\mathrm{s}(W)\cap\mathrm{r}(W)=\emptyset$. Take $h\in C_{c}(\mathrm{s}(W))$ such that $h|_{\mathrm{s}(\mathrm{supp}f)}=1$ and $0\leq h\leq 1$. A direct calculation as in the proof of [21, Lemma 4.3] shows that that $f\ast h=f$ while $h\ast f=0$. Hence, we obtain $\tau(f)=\tau(f\ast h)=\tau(h\ast f)=0$ as required.

### v3

@@ Abstract.
As applications, we obtain the following: 1) a group action is essentially free if every tracial state on the reduced crossed product is canonical and every isotropy group is amenable; 2) if the groupoid $\mathcal{G}$ is second-countable, amenable and essentially free then every (not necessarily faithful) tracial state on the reduced groupoid $C^{*}$-algebra is quasidiagonal.
@@
In recent years, there has been increasing interest in the classification of crossed products of $C^{*}$-algebras $C_{0}(X)\rtimes_{\nu}\Gamma$ arising from discrete amenable group actions on locally compact Hausdorff spaces $\Gamma\curvearrowright X$ (see, e.g., [2, 10, 12, 15, 16, 17, 19, 22, 25, 26, 29, 30]). One of the key ingredients in those proofs is that every tracial state $\tau$ on $C_{0}(X)\rtimes_{\nu}\Gamma$ is canonical in the sense that $\tau=\tau|_{C_{0}(X)}\circ E$, where $E:C_{0}(X)\rtimes_{\nu}\Gamma\to C_{0}(X)$ is the canonical conditional expectation. Actually, it was shown in [14, Theorem 2.7] that every tracial state on the maximal crossed product $C(X)\rtimes\Gamma$ of an action on a compact Hausdorff space $X$ is canonical if and only if the action is essentially free with respect to all invariant probability Radon measures on $X$. On the other hand, every tracial state on the reduced crossed product $C(X)\rtimes_{r}\Gamma$ is canonical if and only if the action of the amenable radical $R_{a}(\Gamma)$ of $\Gamma$ (i.e., the largest amenable normal subgroup in $\Gamma$) on $X$ is essentially free with respect to all invariant probability Radon measures on $X$ (see [34, Corollary 1.12]).
On the other hand, X. Li was able to show that all classifiable $C^{*}$-algebras necessarily arise from twisted étale groupoids (see [20]). Hence, it is natural to consider canonical tracial states on general groupoid $C^{*}$-algebras $C_{\nu}^{*}(\mathcal{G})$ of locally compact Hausdorff étale groupoids $\mathcal{G}$. Similarly, a tracial state $\tau$ on $C_{\nu}^{*}(\mathcal{G})$ is canonical if $\tau=\tau|_{C_{0}(\mathcal{G}^{(0)})}\circ E$, where $E:C_{\nu}^{*}(\mathcal{G})\to C_{0}(\mathcal{G}^{(0)})$ is the canonical conditional expectation. If we let $\mu$ be the uniquely associated invariant probability Radon measure on $\mathcal{G}^{(0)}$ to $\tau|_{C_{0}(\mathcal{G}^{(0)})}$, then it follows from [27, Corollary 1.2] and [28, Corollary 2.4] that if $\mathcal{G}$ is second-countable, then a tracial state $\tau$ on the maximal groupoid $C^{*}$-algebra $C^{*}(\mathcal{G})$ is canonical if and only if $\mathcal{G}$ is essentially free with respect to the associated measure $\mu$ (see Definition 3.3).
@@
where $f_{i}\in C_{0}(X)$ and $\gamma_{i}\in\Gamma$ for $i=1,\dotsc,n$, and $\mathrm{Fix}(\gamma_{i})=\{x\in X:\gamma_{i}x=x\}$ is the set of fixed points of $\gamma_{i}$. The key observation is that $\tau^{\mathrm{Fix}}_{\mu}$ defined on $C_{c}(\mathcal{G})$ is canonical if and only if $\mathcal{G}$ is essentially free with respect to $\mu$ (see Lemma 3.14). Using the fixed point trace, we prove the following main result of this paper which answers Question 1.1:
@@
(Corollary 5.5) Let $\Gamma$ be a discrete group acting on a locally compact Hausdorff space $X$ with an invariant probability Radon measure $\mu$. We consider the following conditions:
@@
Then (1) $\Rightarrow$ (2) [^1] Note that the implication “(2) $\Rightarrow$ (1)” in Corollary B does not hold in general (see Remark 5.6). and (2) + (3) $\Rightarrow$ (1).
If additionally $\Gamma$ is countable and $X$ is second-countable, then (1) $\Leftrightarrow$ (2) + (3).
We end this paper with the following result about quasidiagonal traces, which play a crucial role in the classification of simple nuclear $C^{*}$-algebras (see [36]):
@@
(Corollary 5.16222We also refer the reader to Theorem 5.15 for a more general result.) Let $\mathcal{G}$ be a locally compact, Hausdorff, second-countable, amenable and étale groupoid, which is also essentially free. Then every (not necessarily faithful) tracial state on $C_{r}^{*}(\mathcal{G})$ is quasidiagonal.
@@ 2.1. Basic notions for groupoids
and an inverse map on $\mathcal{G}$ given by $\gamma\mapsto\gamma^{-1}$. These operations satisfy a couple of axioms, including the associativity law and the fact that elements in $\mathcal{G}^{(0)}$ act as units. For $x\in\mathcal{G}^{(0)}$, we define $\mathcal{G}^{x}:=\mathrm{r}^{-1}(x)$ and $\mathcal{G}_{x}:=\mathrm{s}^{-1}(x)$. Moreover, $\mathcal{G}^{x}_{x}=\mathcal{G}^{x}\cap\mathcal{G}_{x}$ is called the isotropy group at $x\in\mathcal{G}^{(0)}$. A subset $Y\subseteq\mathcal{G}^{(0)}$ is called invariant if $\mathrm{r}^{-1}(Y)=\mathrm{s}^{-1}(Y)$, and we define $\mathcal{G}_{Y}:=\mathrm{s}^{-1}(Y)$. For $A,B\subseteq\mathcal{G}$, we define
@@
A locally compact Hausdorff groupoid is a groupoid $\mathcal{G}$ endowed with a locally compact and Hausdorff topology for which the composition, inversion, source and range maps are continuous with respect to the induced topologies.
We say that a locally compact Hausdorff groupoid $\mathcal{G}$ is étale if the range (and hence the source) map is a local homeomorphism, i.e., for any $\gamma\in\mathcal{G}$ there exists an open neighbourhood $U$ of $\gamma$ such that $\mathrm{r}(U)$ is open and $\mathrm{r}|_{U}$ is a homeomorphism. In this case, the fibers $\mathcal{G}_{x}$ and $\mathcal{G}^{x}$ with the induced topologies are discrete for each $x\in\mathcal{G}^{(0)}$, and $\mathcal{G}^{(0)}$ is clopen in $\mathcal{G}$.
A subset $A$ in an étale groupoid $\mathcal{G}$ is called a bisection if the restrictions of $\mathrm{s},\mathrm{r}$ to $A$ are homeomorphisms onto their respective images. It follows from definitions that all open bisections form a basis for the topology of $\mathcal{G}$. As a direct consequence, any function $f\in C_{c}(\mathcal{G})$ can be written as a linear combination of continuous functions whose supports are contained in pre-compact open bisections.
We record the following known result (see, e.g., [32, Lemma 8.4.11]). For convenience of the reader, we provide here a short proof.
@@
Given two open subsets $U$ and $V$ of $\mathcal{G}$, we need to show that $UV$ is also open. Without loss of generality, we can assume that $U$ and $V$ are open bisections with $\mathrm{s}(U)=\mathrm{r}(V)$. Fix $\alpha\in U$ and $\beta\in V$ with $\mathrm{s}(\alpha)=\mathrm{r}(\beta)$. Since the multiplication map is continuous at $(\alpha^{-1},\alpha\beta)$, there exists an open neighbourhood $U_{1}\subseteq U$ of $\alpha$ and an open bisection $W\subseteq\mathrm{r}^{-1}(\mathrm{r}(U_{1}))$ containing $\alpha\beta$ such that $U_{1}^{-1}W\subseteq V$. As $U_{1}\subseteq U$ is also a bisection, we obtain that $W\subseteq UV$. ∎
@@ 2.2. Groupoid $C^{*}$-algebras
Let us now recall different constructions of groupoid $C^{*}$-algebras and their basic properties. Given a locally compact Hausdorff and étale groupoid $\mathcal{G}$, the space $C_{c}(\mathcal{G})$ can be turned into a *-algebra with the following operations: given $f,g\in C_{c}(\mathcal{G})$, we define their convolution and involution by
@@
To tell the difference, we denote the point-wise product by $f\cdot g$.
@@
and the reduced groupoid $C^{*}$-algebra $C^{*}_{r}(\mathcal{G})$ is defined to be the completion of the *-algebra $C_{c}(\mathcal{G})$ with respect to the reduced $C^{*}$-norm $\|\cdot\|_{r}$. It is clear that each left regular representation $\lambda_{x}$ can be extended automatically to a *-homomorphism $\lambda_{x}:C^{*}_{r}(\mathcal{G})\to\mathscr{B}(\ell^{2}(\mathcal{G}_{x}))$.
@@
We will also consider other $C^{*}$-norms between the reduced and the maximal ones. More precisely, we say that a $C^{*}$-norm $\|\cdot\|_{\nu}$ dominates the reduced $C^{*}$-norm if $\|f\|_{\nu}\geq\|f\|_{r}$ for all $f\in C_{c}(\mathcal{G})$. It is worth noticing that $\|\cdot\|_{\nu}\leq\|\cdot\|_{\max}$ always holds. We denote the $C^{*}$-completion of $C_{c}(\mathcal{G})$ with respect to $\|\cdot\|_{\nu}$ by $C^{*}_{\nu}(\mathcal{G})$, called a groupoid $C^{*}$-algebra of $\mathcal{G}$. Similarly, we have a surjective $\ast$-homomorphism
@@
We remark that there is an inclusion map $\iota_{0}:C_{c}(\mathcal{G}^{(0)})\to C_{c}(\mathcal{G})$ given by extending functions by zero on $\mathcal{G}\setminus\mathcal{G}^{(0)}$, and it was recorded in [3, Section 2.2] that $\iota_{0}$ can be extended to an isometric $\ast$-homomorphism $\iota:C_{0}(\mathcal{G}^{(0)})\hookrightarrow C^{*}_{r}(\mathcal{G})$, where the norm on $C_{0}(\mathcal{G}^{(0)})$ is the supremum norm. The same fact holds for any $C^{*}$-norm $\|\cdot\|_{\nu}$ dominating the reduced $C^{*}$-norm. Hence, we will in what follows regard $C_{0}(\mathcal{G}^{(0)})$ as a $C^{*}$-subalgebra in $C^{*}_{\nu}(\mathcal{G})$ without further explanation.
From [31, Proposition II.4.2] (see also [4, Section 2.2]) we have that any element of $C^{*}_{r}(\mathcal{G})$ can be regarded as a $C_{0}$-function on the groupoid $\mathcal{G}$. Indeed, there exists a linear and contractive map $j:C^{*}_{r}(\mathcal{G})\to C_{0}(\mathcal{G})$ given by
@@
for $a\in C^{*}_{r}(\mathcal{G})$ and $\gamma\in\mathcal{G}$. On $C_{c}(\mathcal{G})\cup C_{0}(\mathcal{G}^{(0)})$ the map $j$ is nothing but the identity map. The reduced groupoid $C^{*}$-algebra also admits a faithful conditional expectation $E:C^{*}_{r}(\mathcal{G})\to C_{0}(\mathcal{G}^{(0)})$ defined by
@@
We end this subsection with an elementary fact, and leave its proof (which is a relatively straightforward computation) to the reader.
@@ 2.3. Tracial states
If $\tau$ is a tracial state on the groupoid $C^{*}$-algebra $C^{*}_{\nu}(\mathcal{G})$, then $\tau|_{C_{0}(\mathcal{G}^{(0)})}$ is a state on $C_{0}(\mathcal{G}^{(0)})$, which corresponds to a (positive) probability Radon measure $\mu$ on $\mathcal{G}^{(0)}$ according to the Riesz representation theorem. In other words, we have
(2.6) $\tau(f)=\int_{\mathcal{G}^{(0)}}f\mathrm{d}\mu\quad\text{for any}\quad f\in C_{0}(\mathcal{G}^{(0)}).$
We call $\mu$ the measure associated to $\tau$ and we also denote this measure by $\mu_{\tau}$. It is actually invariant in the following sense (see, e.g., [21, Lemma 4.1]):
@@
(2.7) $\alpha_{B}:\mathrm{s}(B)\to\mathrm{r}(B)\quad\text{given by}\quad x\mapsto\mathrm{r}((\mathrm{s}|_{B})^{-1}(x))\quad\text{for}\quad x\in\mathrm{s}(B).$
A Borel measure $\mu$ on $\mathcal{G}^{(0)}$ is called invariant (cf. [31, Definition I.3.12]) if for any open bisection $B$ in $\mathcal{G}$, we have $\mu|_{\mathrm{r}(B)}=(\alpha_{B})_{\ast}(\mu|_{\mathrm{s}(B)})$. For an invariant measure $\mu$ on $\mathcal{G}^{(0)}$, its support $\mathrm{supp}\mu$ is an invariant subset of $\mathcal{G}^{(0)}$. We also define
@@
Conversely, let $\mu$ be an invariant probability Radon measure on $\mathcal{G}^{(0)}$. Then it follows from [21, Lemma 4.2] (see also [31, Proposition II.5.4]) that
@@
is a tracial state on $C^{*}_{r}(\mathcal{G})$, called the tracial state associated to $\mu$. Similarly, we can also consider the tracial state associated to $\mu$ on $C^{*}_{\nu}(\mathcal{G})$ (with the same notation)
@@
Let $\mathcal{G}$ be a locally compact Hausdorff and étale groupoid, and $\|\cdot\|_{\nu}$ be a $C^{*}$-norm on $C_{c}(\mathcal{G})$ dominating the reduced $C^{*}$-norm. A tracial state $\tau$ on $C^{*}_{\nu}(\mathcal{G})$ is called canonical if $\tau=\tau_{\mu}$, where $\mu$ is the measure associated to $\tau$ on $\mathcal{G}^{(0)}$.
The following is elementary but useful in what follows:
@@
A tracial state $\tau$ on $C^{*}_{\nu}(\mathcal{G})$ is canonical if and only if $\tau(f)=0$ for all $f\in C_{c}(\mathcal{G}\setminus\mathcal{G}^{(0)})$.
@@
The forward implication is clear. Now we assume that $\tau(f)=0$ for all $f\in C_{c}(\mathcal{G}\setminus\mathcal{G}^{(0)})$. As $C_{c}(\mathcal{G})$ is dense in $C^{*}_{\nu}(\mathcal{G})$, it suffices to show that $\tau(f)=\tau_{\mu}(f)$ for any $f\in C_{c}(\mathcal{G})$, where $\mu$ is the measure associated to $\tau$. Since $\mathcal{G}$ is étale, we have the decomposition $f=f|_{\mathcal{G}^{(0)}}+f|_{\mathcal{G}\setminus\mathcal{G}^{(0)}}$ in $C_{c}(\mathcal{G})$. Hence by assumption, we have $\tau(f)=\tau(f|_{\mathcal{G}^{(0)}})$, which finishes the proof by the definition of the canonical trace (2.6). ∎
It is interesting to know when every tracial state on $C^{*}_{\nu}(\mathcal{G})$ is canonical. Recall from [21, Lemma 4.3] (see also [31, Proposition II.5.4]) that if the groupoid $\mathcal{G}$ is principal, then every tracial state on $C^{*}_{r}(\mathcal{G})$ is canonical. We would like to weaken the condition of being principal to the following notion of essential freeness.
@@
The following proposition is perhaps known to experts at least for the maximal $C^{*}$-norm (see [27, Corollary 1.2]). We provide here a self-contained proof, because we cannot find the explicit statement we need in the literature.
@@
Our proof here is mainly inspired by the one for [18, Proposition 1.1]. By Lemma 3.2, it suffices to show that $\tau(f)=0$ for any $f\in C_{c}(\mathcal{G}\setminus\mathcal{G}^{(0)})$. By decomposing $f$ into its positive part and negative part, it suffices to show that $\tau(f)=0$ for any $f\in C_{c}(\mathcal{G}\setminus\mathcal{G}^{(0)})$ which is point-wise non-negative. Using an argument of partitions of unity, we can additionally assume that $\mathrm{supp}f\subseteq B$ for some pre-compact open bisection $B$ such that $\overline{B}\subseteq B_{0}$ for another pre-compact open bisection $B_{0}\subseteq\mathcal{G}\setminus\mathcal{G}^{(0)}$.
First we suppose that $\mathrm{supp}f\cap\{\gamma\in B:\mathrm{s}(\gamma)=\mathrm{r}(\gamma)\}=\emptyset$. Then for any $\gamma\in\mathrm{supp}f$, there exists an open neighbourhood $W_{\gamma}$ of $\gamma$ such that $\mathrm{s}(W_{\gamma})\cap\mathrm{r}(W_{\gamma})=\emptyset$. Since $\mathrm{supp}f$ is compact, we can choose a finite cover $\{W_{\gamma_{1}},\dotsc,W_{\gamma_{N}}\}$ for $\mathrm{supp}f$, and take a partition of unity $\{\rho_{\gamma_{1}},\dotsc,\rho_{\gamma_{N}}\}$. Then we can write $f=\sum_{n=1}^{N}(\rho_{n}\cdot f)$, where $\rho_{n}\cdot f$ denotes the point-wise product of $\rho_{n}$ and $f$ as in Section 2.2. For each $n=1,\dotsc,N$, take $h_{n}\in C_{c}(\mathrm{s}(W_{\gamma_{n}}))$ such that $h_{n}|_{\mathrm{s}(\mathrm{supp}(\rho_{n}\cdot f))}=1$ and $0\leq h_{n}\leq 1$. A direct calculation as in the proof of [21, Lemma 4.3] shows that that $(\rho_{n}\cdot f)\ast h_{n}=\rho_{n}\cdot f$ while $h_{n}\ast(\rho_{n}\cdot f)=0$. Hence, we obtain $\tau(\rho_{n}\cdot f)=\tau((\rho_{n}\cdot f)\ast h_{n})=\tau(h_{n}\ast(\rho_{n}\cdot f))=0$, which implies that $\tau(f)=\sum_{n=1}^{N}\tau(\rho_{n}\cdot f)=0$, as required.